# Cell divisions imprint long lasting elastic strain fields in epithelial tissues

### v1

@@ II.2 Viscous dissipation determines tissue relaxation dynamics
Our analysis of the strain in the ablated tissue after it has relaxed allowed us to study its elastic properties. Now, we explore dynamics of tissue strain as it approaches the equilibrium, which will allow us to study dissipative processes in fly wing epithelium. We consider two possible dissipation mechanisms: viscosity and friction with a substrate.
@@
Interestingly, in a continuum model of a tissue as an elastic sheet, these two dissipation mechanisms lead to qualitatively different relaxation dynamics upon insertion of a force dipole. In particular, in an elastic sheet with only viscous dissipation, all strain field components decays as $1/r^{2}$ with the distance from the force dipole throughout the relaxation with the amplitude evolving in time, see SI and Fig. 2 (A.i). However, in an elastic sheet with only frictional dissipation, the strain field propagates through the sheet: at distances smaller than the propagation front, the strain field converges to the steady state solution and rapidly decays beyond the front, see SI and Fig. 2 (A.ii).
To quantify the dynamics of tissue relaxation following the laser ablation, we measure the strain field in time intervals of $4.5s$, starting from $1s$ after the ablation, (SI Fig. 7). We find that the angular strain pattern is the same as in the mechanical equilibrium one, see Movie 1. We analyse the dynamics of the strain radial profile by calculating the second angular Fourier mode of the strain field at all measured timepoints, see Fig. 2 (B.i) and SI Fig. 7. We find that at all timepoints the strain field is consistent with $1/r^{2}$ decay and we do not see any signature of a propagating front, see Fig. 2. Therefore, we conclude that the viscosity is the dominant mode of dissipation in the fly wing epithelium on timescales above $1s$, corresponding to the initial timepoint we observe after the ablation, and below $T_{e}\approx 50s$.
@@
To make sure that a predicted propagation front could be observed in presence of frictional dissipation we performed vertex model simulations of laser ablation that include only the frictional dissipation, see SI for details. We indeed find a clear propagation of the strain field that converges to the steady state value as the propagation front expands (2 (B.ii) and SI Fig. 8), as predicted by the analystical calculation (2 (A.ii)).
@@
Figure 2: Time evolution the second Fourier mode $m=2$ of the strain field $xy$ component induced by a force dipole for: (A.i) linear elastic sheet with viscous dissipation, (A.ii) linear elastic sheet with frictional dissipation, (B.i) laser ablation experiment (B.ii) vertex model simulation with frictional dissipation. Dashed line indicates the power-law $r^{-2}$ decay for reference.
Table 1: Parameters obtained by fitting the continuum theory to spatio-temporal pattern of strain field accumulated in the tissue after the ablation.
exp. $d_{0}~{}[\mu m]^{2}$ $\tilde{d}_{xx}~{}[\mu m]^{2}$ $\tilde{d}_{xy}~{}[\mu m]^{2}$ $\mu$ $\tau_{1}~{}(s)$ $\tau_{2}~{}(s)$
1 $16\pm 1$ $15\pm 1$ $2\pm 1$ $1.9\pm 0.1$ $11\pm 1$ $12\pm 1$
2 $15\pm 2$ $31\pm 2$ $0\pm 2$ $2.1\pm 0.3$ $48\pm 2$ $55\pm 2$
@@
Our continuum model of an elastic sheet provides the dynamical strain propagator ${G}_{ij}(\bm{r},t)={G}_{ij}(\bm{r}){T}_{ij}(t)$, where no sum over repeated indices is performed. ${G}_{ij}(\bm{r})$ is the steady state propagator, and ${T}_{ij}(t)$ is the relaxation matrix whose components contains two relaxation factors $g_{1}(t)=1-\exp{(-t/\tau_{1})}$ and $g_{2}(t)=1-\exp{(-t/\tau_{2})}$, where the two relaxation timescales
@@
are related to the bulk and shear elastic constants, and bulk and shear viscosities $\overline{K},2K,\overline{\eta},\text{ and }2\eta$, respectively (see SI). To extract $\tau_{1}$ and $\tau_{2}$ from experimental data, as described in SI and shown in SI Fig. 1. We find that the two relaxation timescales are very similar, although they vary between experiments, see Table 1. It is at the moment unclear whether the difference between the relaxation timescales in two experiments stems from variability of elastic constants or viscosity, or some other factor. While this will be an important open question to resolve in future, further investigation of relaxation timescale variability is beyond the scope of this work.
So far, our detailed analysis of the strain in the fly wing epithelium in response to the laser ablation shows that the tissue behaves as a two-dimensional viscoelastic solid. With this knowledge in hand, we can now infer force dipoles generated by cell divisions, and thereby study their mechanics, by simply observing the dynamics of the surrounding tissue strain caused by the division, without the need to directly perturb the tissue.
@@ III Elastic deformation imprinted by a cell division
Cell division is a complex process during which the mother cell is split into two daughter cells. Here, we aim to characterise the mechanical forces generated by the dividing cell by analysing the strain field a dividing cell induces in the fly wing epithelium.
@@
We analysed cell divisions in previously published experiments Dye et al. (2021) on explanted D. melanogaster wing imaginal discs  [^2] Note that this developmental stage preceeds the pupal stage in which linear laser ablations were performed Piscitello-Gómez et al. (2023) and although both are fruit fly wing tissues, it is possible that tissue elastic constants have changed between the two stages.. Wing discs were cultured over about $13$ hours, during which the tissue was imaged every $5$ minutes, and individual cells were segmented and tracked over time. The progress of a typical cell division is shown in Fig. 3 (A). Before the actual cell division, the mother appears to inflate Fig. 3 (A.ii), which is not a real increase in cell size, but is rather a consequence of the cell nucleus arriving to the apical surface of the cell to divide Gibson et al. (2011); Strzyz et al. (2016). This transient inflation is followed by the creation of the new cell boundary, and the two daughter cells become visible Fig. 3 (A.iii). For each cell division, we define the time of the new bond creation as the reference cell division time $t=0$.
@@
Fig. 3 (A.iv) qualitatively summarizes these experimental observations. For each cell division, we quantify the tissue strain field in the vicinity of the dividing cell, as described in Section II, relative to the starting timepoint $t_{0}=-60~{}\rm{min}$ before the cell division time. We show an example of the cell strain field induced by a dividing cell in Fig. 3 (B). The timescale on which the cell division forces evolve, which is on the order of tens of minutes, is much larger than the relaxation timescale extracted in Section II, which is below one minute. Therefore, the observed strain field dynamics is largely due to changes of the force dipole generated by the cell division, and not due to much faster stress propagation through the tissue.
@@
Figure 3: (A) Snapshots of a cell division in three stages. (A.i) As the mother cell starts the division process, (A.ii) it rounds up, corresponding to the inflation, then (A.iii) it shrinks and divides into the daughter cells. The red dots show the mother and the daughter cells. (A.iv) A schematic representation of the cell area of the mother cell (brown) and the daughter cells (orange) during the division process. The symbols highlight the stage of the dividing cell before the division, at the peak of inflation, and after the division. The new bond forms at $t=0$, and $t_{0}$ is the reference state. (B) Strain $\tilde{U}_{xy}$ of the cells near a typical cell division at the three stages shown in (A) with $t_{0}=-60~{}[\rm{min}]$. (C) Ensemble average ($N=682$) of $\tilde{U}_{xy}$ around the dividing cell: (C.i) at the peak of the inflation $t=-20~{}[\rm{min}]$ (corresponding to (A.ii) and (B.ii)), and (C.ii) after the division $t=30~{}[\rm{min}]$ (corresponding to (A.iii) and (B.iii)). (D) The force dipole generated by a cell division as a function of time. (E) A schematic summary of the force dipoles during stages of a cell division.
@@
We find that a single cell division generates weaker forces compared to those induced by laser ablation (compare Fig. 1 A and Fig. 3 B), and the corresponding strain field is more susceptible to noise. Therefore, we average the strain field from multiple cell divisions ($N=682$). We first align the centers of the cell division, defined as the area weighted average of daughter cell centers at the cell division time ($t=0$). We next rotate each division such that the orientation of the line connecting daughter cell centers is oriented along the $x-$axis at the cell division time. We show the resulting strain field component $\tilde{U}_{xy}$ in Fig. 3 (C) at times $t=-20~{}\rm{min}$, corresponding to the peak of the cell inflation, and $t=30~{}\rm{min}$ after the cell division. In Fig. 3 (C.i), the 2-fold symmetry of the angular strain pattern indicates that the isotropic force dipole is dominant, consistent with the isotropic cell inflation. In Fig. 3 (C.ii), the pattern is closer to 4-fold symmetric one, such as would be induced by an anisotropic force dipole.
@@
We fit the linear elastic theory Eq. 2 to the experimentally measured strain field at each time point and extract the resulting force dipole $d_{ij}(t)$. We find that the isotropic force dipole component $d_{0}$ rises sharply during the mother cell inflation and peaks at about $t=-20~{}\rm{min}$ (magenta line in Fig. 3 D). However, by the time of the new bond formation, the isotropic component has disappeared and remains negligible at later times. This result shows that cell does not noticeably grow during the division, but rather that the division reorganises the available material into the two daughter cells. The traceless part of the force dipole $\tilde{d}_{xx}$ grows as the new bond is formed and persists in the tissue multiple hours after the division (green line in Fig. 3 D), indicating that the division-induced strains remain imprinted in the tissue long after the division has been executed. Over time, $\tilde{d}_{xx}$ slowly decays while its fluctuations increase, representing the effect of other cell divisions and rearrangements that appear in the tissue. We estimate the tissue fluidization time from the decay of $\tilde{d}_{xx}$ in time by fitting an exponential function to it. We find the fluidization timescale to be $\tau_{\alpha}\simeq 3.5\pm 0.7$ minutes, over which $\tilde{d}_{xx}$ vanishes and, therefore, the strain field induced by the division is erased. Note that component $\tilde{d}_{xy}$ remains negligible at all times.
@@
Finally, we note that at the time of the peak of $d_{0}$, there is a small but clear negative peak in $\tilde{d}_{xx}$. To interpret this observation, we have to remember that our strain measurements are relative to the state of the cells at $t_{0}=-60~{}\rm{min}$. Since the mother cell attains an almost perfectly circular shape during the inflation, any initial anisotropy of the mother cell at time $t_{0}$ gives rise to an anisotropic force dipole component during the inflation. Therefore, the observed negative value of $\tilde{d}_{xx}$ indicates that mother cells are slightly elongated along the future division axis, consistent with the well known Hertwig’s rule Hertwig (1884); Théry et al. (2007); Gibson et al. (2011); Lisica et al. (2022); Middelkoop et al. (2023). We summarise these results in Fig. 3 (E).
@@ IV Discussion
In this work, we devised a full spatio-temporal analysis of tissue strain induced by two types of force dipole fields: linear laser ablations and cell divisions. In this way, we have shown that the developing Drosophila wing epithelium behaves as a two-dimensional elastic solid with viscous dissipation. We also measured its two-dimensional Poisson ratio, which we find to be negative, consistent with the values reported in the literature Dye et al. (2021); Piscitello-Gómez et al. (2023).
We characterized the dynamics of mechanical forces generated by a dividing cell throughout the cell division process. We found that strain fields imprinted by a cell division are slowly erased on the timescale of $\tau_{\alpha}=3.5\pm 0.7~{}\rm{hr}$, which can be understood as a fluidization timescale of the Drosophila wing tissue. Furthermore, our analysis shows that there is no cell growth during a division and stresses induced by the cell division correspond to those of a pure shear force dipole. Finally, we found that Hertwig’s rule applies in the D. melanogaster wing disc Gibson et al. (2011); Mao et al. (2013).
An important implication of our work is that the Drosophila wing epithelium is a solid whose fluidization time $\tau_{\alpha}$ is not negligible on the developmental timescales. For example, large scale shape changes that the wing undergoes during eversion are executed on a similar timescale Fuhrmann et al. (2023). A possible physical picture of the Drosophila wing epithelium that could account for our observations, is that of a glassy material, or as was put forward in glass literature, a solid that flows Dyre (2023); Ozawa and Biroli (2023); Tahaei et al. (2023). In other words, the long-range elastic interactions between cells would lead to correlated glassy dynamics, which has been reported in tissue cultures Angelini et al. (2011); Park et al. (2015), and it remains to be seen whether they indeed play a role in developing tissues. In summary, our results quantify force dipoles generated by dividing cells in developing D. melanogaster wing epithelia, as well es elastic interactions between cells, which paves the way to development tissue mechanics theory and motivates further experimental study of collective cell dynamics.

### v2

@@ II.2 Viscous dissipation determines tissue relaxation dynamics
Our analysis of the strain in the ablated tissue after it has relaxed allows us to study its elastic properties. Now, we explore the transient dynamics of tissue strain as it approaches the mechanical equilibrium, which will allow us to study dissipative processes in fly wing epithelium. We consider two possible dissipation mechanisms: viscosity and friction with a substrate.
@@
Interestingly, in a continuum model of tissue as an elastic sheet, these two dissipation mechanisms lead to qualitatively different relaxation dynamics upon insertion of a force dipole. In particular, in an elastic sheet with only viscous dissipation, all strain field components decay as $1/r^{2}$ with the distance from the force dipole throughout the relaxation with the amplitude evolving in time, see SI and Fig. 2 (A.i). However, in an elastic sheet with only frictional dissipation, the strain field propagates through the sheet: at distances within the propagation front, the strain field converges to the steady state solution and rapidly decays beyond the front, see SI and Fig. 2 (A.ii).
@@
To compare the strain field dynamics in viscous and frictional dissipation models within a cellular system, we performed vertex model simulations of laser ablation that include only the frictional dissipation and compare it with a viscous dissipation, see SI for details. As discussed above, the signature difference between viscous and frictional dissipation is the power law decay of the shape $1/r^{2}$ in the former and the existence of a propagation front in the latter. With a frictional dissipation, we indeed find a clear propagation of the strain field that converges to the steady state value as the propagation front expands (2 (B.ii) and SI Fig. 12 (B)), consistent with the analytical calculation (2 (A.ii)). In contrast, with viscous dissipation, we observe that the strain field decays as $1/r^{2}$ in all time points (2 (B.i) and SI Fig. 12 (A)).
To quantify the dynamics of tissue relaxation following the laser ablation, we measure the strain field in time intervals of $4.5s$, starting from $1s$ after the ablation, (SI Fig. 8). We find that the angular strain pattern is the same as in the mechanical equilibrium one, see Movie 1. We analyze the dynamics of the strain radial profile by calculating the second angular Fourier mode of the strain field at all measured timepoints, see Fig. 2 (C.i) and SI Fig. 8. We find that the strain field is consistent with $1/r^{2}$ decay and we do not see any signature of a propagating front, see Fig. 2. Therefore, we conclude that the viscosity is the dominant mode of dissipation in the fly wing epithelium on timescales below $T_{e}\approx 30s$.
@@
Figure 2: Time evolution the second Fourier mode $m=2$ of the strain field $xy$ component induced by a force dipole for: (A.i) linear elastic sheet with viscous dissipation, (A.ii) linear elastic sheet with frictional dissipation, vertex model simulation with (B.i) viscous dissipation and, (B.ii) frictional dissipation. (C) Laser ablation experiment. (C.i) Time evolution of $\tilde{U}_{xy}$ shows a radial profile as $r^{-2}$, consistent with a viscous dissipation mechanism. (C.ii) Quantifying the viscous relaxation of strain as described in the text, which allows the measurement of viscous relaxation timescales. The black dashed line indicates the power-law $r^{-2}$ decay for reference. Each data point represents the strain field average over $5$ smaller bins and the errors are the corresponding standard deviations over those bins.
@@
Our continuum model of an elastic sheet provides the dynamical strain propagator ${G}_{ij}(\bm{r},t)={G}_{ij}(\bm{r}){T}_{ij}(t)$, where no sum over repeated indices is performed. ${G}_{ij}(\bm{r})$ is the steady state propagator, and ${T}_{ij}(t)$ is the relaxation matrix whose components contain two relaxation factors $g_{1}(t)=1-\exp{(-t/\tau_{1})}$ and $g_{2}(t)=1-\exp{(-t/\tau_{2})}$, where the two relaxation timescales
@@
are related to the bulk and shear elastic constants, and bulk and shear viscosities $\overline{K},2K,\overline{\eta},\text{ and }2\eta$, respectively (see SI I.B.(ii)). We extract $\tau_{1}$ and $\tau_{2}$ from experimental data, see Methods, SI, and SI Fig. 13. We find that the two relaxation timescales are very similar: $\tau_{1}=11\pm 1s$ and $\tau_{2}=12\pm 1s$, where intervals of uncertainty correspond to the uncertainty of the fit, see Methods.
As mentioned above, the Experiment 1 presented here was selected based on our ability to segment and track cells in a radius of $40~{}\mu m$ surrounding the laser ablation, which was not possible for most of the other experiments. Here, we highlight that in Experiment 2, the angular and radial pattern of the strain field are consistent with that of an elastic sheet with viscous dissipation on times comparable to the relaxation timescales $\tau_{1}$ and $\tau_{2}$ extracted above, as reported in the SI Fig. 13. Interestingly, on longer times radial profile of strain departs from $r^{-2}$ and continues to evolve throughout the experiment. Similar additional motion is observed even in the Experiment 1 reported above, but only after about $2~{}\rm{min}$, see SI Fig. 11. This suggests that additional dynamics are due to a wound-healing response in the tissue and understanding this behavior would require a dedicated experimental exploration in the future.
In summary, our detailed analysis of the strain in the fly wing epithelium in response to the laser ablation shows that the tissue behaves as a two-dimensional viscoelastic solid. These results suggest that we can now infer force dipoles generated by cell divisions, and thereby study their mechanics, by simply observing the dynamics of the surrounding tissue strain caused by the division, without the need to directly perturb the tissue.
@@ III Elastic deformation imprinted by a cell division
Cell division is a complex process during which the mother cell is split into two daughter cells. Here, we aim to characterize the mechanical forces generated by the dividing cell from the analysis of the strain field that the dividing cell induces in the epithelium. We analyze cell divisions in previously published experiments Dye et al. (2021) on explanted D. melanogaster wing imaginal discs. We are interested in the generic features of the response of epithelial tissues to a mechanical perturbation, such as the laser ablation or a cell division. Therefore, we aim to extract normalised force dipoles in the surface plane generated by cell divisions using the same approach as with the laser ablation experiments. However, we perform the analysis independently of any results obtained from laser ablations in the pupal wing.
The wing disc developmental stage precedes the pupal stage in which linear laser ablations analysed above were performed Piscitello-Gómez et al. (2023), and their mechanical moduli may be different. However, a quantification of the strain field generated by cell divisions is more convenient in the wing disc stage compared to the pupal wing stage. Cell divisions in the pupal wing differ from divisions in the wing disc in that they are patterned in time and space, with most cells dividing at least once within several hours and the tissue exhibits strong tissue shear flows and active cell rearrangements Etournay et al. (2016). Therefore, analysis of the strain field induced by individual divisions is practically impossible in the pupal wing due to additional dynamics in the tissue. However, in both tissues, the apical surfaces are mechanically special as they contain a high concentration of actomyosin cortex and adherens junctions between cells that are important for morphogenesis. Therefore, it is likely that apical mechanics plays a significant role in the overall tissue mechanics in both tissues.
@@
Wing discs were cultured over about $13$ hours, during which the tissue was imaged every $5$ minutes, and individual cells were segmented and tracked over time. The progress of a typical cell division is shown in Fig. 3 (A). Before the actual cell division, the mother appears to inflate Fig. 3 (A.ii). This is not a real increase in cell size, but is rather a consequence of the cell nucleus arriving at the apical surface of the cell to divide and transiently stretching the cell apical surface Gibson et al. (2011); Strzyz et al. (2016). This transient inflation is followed by the creation of the new cell boundary, and the two daughter cells become visible Fig. 3 (A.iii). For each cell division, we define the time of the new bond creation as the reference cell division time $t=0$.
@@
Fig. 3 (A.iv) qualitatively summarizes the experimental observations. For each cell division, we quantify the tissue strain field in the vicinity of the dividing cell, as described in Section II, relative to the starting timepoint $t_{0}=-60~{}\rm{min}$ before the cell division time. We show an example of the cell strain field induced by a dividing cell in Fig. 3 (B). The timescale on which the cell division forces evolve, which is on the order of tens of minutes, is much larger than the minutes relaxation timescale reported in the wing disc Dye et al. (2021). Therefore, the strain field dynamics observed over tens of minutes is largely due to changes of the force dipole generated by the cell division, and not due to much faster stress propagation through the tissue. Although it is possible that on these longer timescales cells adapt to the experienced stress and adapt their reference shape, a clear positive correlation between stress and strain has been reported in our past work Dye et al. (2021). Therefore, it is reasonable to interpret the change in cell elongation as a measure of elastic strain, even if the adaptation to stress effectively renormalizes the elastic constants, as discussed in Ref. Dye et al. (2021).
@@
Figure 3: (A) Snapshots of a cell division in three stages. (A.i) As the mother cell starts the division process, (A.ii) it rounds up, corresponding to the inflation, and then (A.iii) it shrinks and divides into the daughter cells. The red dots show the mother and the daughter cells. (A.iv) A schematic representation of the cell area of the mother cell (brown) and the daughter cells (orange) during the division process. The symbols highlight the stage of the dividing cell before the division, at the peak of inflation, and after the division. The new bond forms at $t=0$, and $t_{0}$ is the reference state. (B) Strain $\tilde{U}_{xy}$ of the cells near a typical cell division at the three stages shown in (A) with $t_{0}=-60~{}[\rm{min}]$. (C) Ensemble average ($N=682$) of $\tilde{U}_{xy}$ around the dividing cell: (C.i) at the peak of the inflation $t=-20~{}[\rm{min}]$ (corresponding to (A.ii) and (B.ii)), and (C.ii) after the division $t=30~{}[\rm{min}]$ (corresponding to (A.iii) and (B.iii)). The radial extent of the plots is $6\mu m$. (D) The angular Fourier transform of the $\tilde{U}_{xy}$ at a distance of $r=3~{}\mu m$ from the center of the mother cell at $t=-20~{}[\rm{min}]$ (D.i) and at $t=30~{}[\rm{min}]$ (D.ii).
Figure 4: (A) The fit of linear elastic theory to the measured strain field around cell divisions at (A.i) the peak of inflation $t=-20~{}[\rm{min}]$ and (A.ii) after division $t=30~{}[\rm{min}]$. Each data point shows the average strain field in polar bins and the error bars are the standard deviation over these bins. (B) The force dipole generated by a cell division as a function of time. (C) A schematic summary of the force dipoles during stages of a cell division.
@@
We find that a single cell division generates weaker forces compared to those induced by laser ablation (compare Fig. 1 A and Fig. 3 B), and the corresponding strain field is more susceptible to noise. Therefore, we average the strain field from multiple cell divisions ($N=682$). We first align the centers of the cell division, defined as the area-weighted average of daughter cell centers at the cell division time ($t=0$). We next rotate each division such that the orientation of the line connecting daughter cell centers is oriented along the $x-$axis at the cell division time. We show the resulting strain field component $\tilde{U}_{xy}$ in Fig. 3 (C) and the corresponding angular Fourier modes (D) at times $t=-20~{}\rm{min}$, corresponding to the peak of the cell inflation, and $t=30~{}\rm{min}$ after the cell division. In Fig. 3 (C.i), the 2-fold symmetry of the angular strain pattern indicates that the isotropic force dipole is dominant, consistent with the isotropic cell inflation. In Fig. 3 (C.ii), the pattern is closer to a 4-fold symmetric one, such as would be induced by an anisotropic force dipole. Finally, we find that radial profiles of the Fourier modes are consistent with $r^{-2}$ decay (SI Fig. 15), as was the case with laser ablation experiments above.
@@
We fit the linear elastic theory Eq. 2 to the experimentally measured strain field at each time point and extract the resulting force dipole $d_{ij}(t)$. Comparison between the fit and the data is shown for the $xy$ component of the strain field at $t=-20\rm{min}$ and $t=30\rm{min}$ in Fig. 4 B. We find that the isotropic force dipole component $d_{0}$ rises sharply during the mother cell inflation and peaks at about $t=-20~{}\rm{min}$ (magenta line in Fig. 4 B). However, by the time of the new bond formation, the isotropic component has disappeared and remains negligible at later times. This result shows that cells do not noticeably grow during the division, but rather that the division reorganizes the available material into the two daughter cells. The traceless part of the force dipole $\tilde{d}_{xx}$ grows as the new bond is formed and persists in the tissue multiple hours after the division (green line in Fig. 4 B), indicating that the division-induced strains remain imprinted in the tissue long after the division has been executed. Over time, $\tilde{d}_{xx}$ slowly decays while its fluctuations increase, representing the effect of other cell divisions and rearrangements that appear in the tissue. We estimate the tissue fluidization time from the decay of $\tilde{d}_{xx}$ in time by fitting an exponential function to it. We find the fluidization timescale to be $\tau_{\alpha}\simeq 3.5\pm 0.7$ hours (see Methods), over which $\tilde{d}_{xx}$ vanishes and, therefore, the strain field induced by the division is erased. Note that component $\tilde{d}_{xy}$ remains negligible at all times.
@@
Finally, we note that at the time of the peak of $d_{0}$, there is a small but clear negative peak in $\tilde{d}_{xx}$. To interpret this observation, we have to remember that our strain measurements are relative to the state of the cells at $t_{0}=-60~{}\rm{min}$. Since the mother cell attains an almost perfectly circular shape during the inflation, any initial anisotropy of the mother cell at time $t_{0}$ gives rise to an anisotropic force dipole component during the inflation. Therefore, the observed negative value of $\tilde{d}_{xx}$ indicates that mother cells are elongated along the future division axis, consistent with the well-known Hertwig’s rule Hertwig (1884); Théry et al. (2007); Gibson et al. (2011); Lisica et al. (2022); Middelkoop et al. (2023). We summarise these results in Fig. 4 (C).
@@ IV Discussion
In this work, we devised a full spatio-temporal analysis of tissue strain induced by two types of force dipole fields: linear laser ablations and cell divisions. In this way, we have shown that the developing Drosophila wing epithelium can be described as a two-dimensional elastic solid with viscous dissipation. We also measured its two-dimensional Poisson ratio, which we find to be negative, consistent with the values reported in the literature Dye et al. (2021); Piscitello-Gómez et al. (2023). Our analysis was based on laser ablation of the wing tissue during pupal development, at which time the wing consists of two apposed epithelial layers. Although the experimental observations are consistent with a model of the wing apical surface as a linear elastic sheet, in the future it might be interesting to probe the three-dimensional structure response of the tissue with deeper ablations and in wings where the two epithelial layers have been separated Sun et al. (2021).
We characterized the dynamics of mechanical forces generated by a dividing cell throughout the cell division process. We found that the strain fields imprinted by a cell division are slowly erased on the timescale of $\tau_{\alpha}=3.5\pm 0.7~{}\rm{hr}$, which can be understood as a fluidization timescale of the Drosophila wing tissue. Furthermore, our analysis shows that there is no cell growth during a division, and stresses induced by the cell division correspond to those of a pure shear force dipole. Finally, we found that Hertwig’s rule applies in the D. melanogaster wing disc Gibson et al. (2011); Mao et al. (2013).
An important implication of our work is that the Drosophila wing epithelium is a solid whose fluidization time $\tau_{\alpha}$ is not negligible on the developmental timescales. For example, large-scale shape changes that the wing undergoes during eversion are executed on a similar timescale Fuhrmann et al. (2023). A possible physical picture of the Drosophila wing epithelium that could account for our observations is that of a glassy material, or as was put forward in the glass literature, a solid that flows Dyre (2023); Ozawa and Biroli (2023); Tahaei et al. (2023). In other words, the long-range elastic interactions between cells would lead to correlated glassy dynamics, which has been reported in tissue cultures Angelini et al. (2011); Park et al. (2015), and it remains to be seen whether they indeed play a role in developing tissues. In summary, our results quantify force dipoles generated by dividing cells in developing D. melanogaster wing epithelia, as well as elastic interactions between cells, which paves the way to the development of tissue mechanics theory and motivates further experimental study of collective cell dynamics.